# Rethinking the effects of data contamination in Code Intelligence

### v2

@@ VII-A Code Intelligence Tasks on PLMs and LLMs
Recent years have witnessed extensive research on code intelligence tasks based on PLMs/LLMs, such as code generation [47, 48], code summarization [49, 50], and code translation [51, 38]. Early efforts, such as CodeBERT [52] and GraphCodeBERT [8], jointly modeled NL and PL, enhancing semantic understanding of code through contrastive learning and structural awareness. Subsequently, PLBART [11], based on an encoder-decoder architecture, demonstrated strong transfer capabilities. CodeT5 [12] further pre-trained on T5 [53], achieving state-of-the-art performance on a variety of code understanding and generation tasks. On the LLM front, Codex [15] and CodeGen [17] have shown remarkable capabilities in multi-turn reasoning and complex programming scenarios. More recent works, such as StarCoder [18] and WizardCoder [54] explore the upper performance bounds of open-source LLMs in code intelligence, emphasizing the importance of model scale, instruction tuning, and multilingual training for improving code comprehension and generation abilities.
VII-B Studies On Data Contamination
Data contamination studies have been investigated for years, many are dedicated to contamination detection. For instance, Golchin et al. [25] proposed both instance-level and partition-level contamination detection, while Dong et al. [55] proposed CDD, using output distributions to detect contamination in LLMs on both code and NLP tasks. Deng et al. [56], instead, proposed TS-Guessing to detect contamination by examining whether LLMs can produce masked wrong answers.
@@ VII-B Studies On Data Contamination
As for research exploring contamination effects, early research was conducted in the NLP field. For example, Magar et al. [24] explored to what extent BERT exploits or just memorises the contaminated data for downstream tasks. Jiang et al. [45] explored the input-only/output-only contamination effects on GPT-2. Subsequently, contamination has also been studied in the code intelligence field, Riddell et al. [23] and Ramos et al. [22] quantified the contamination effects on LLM-based code generation and bug fixing, respectively. Cao et al. [21] extracted data before and after LLMs’ cut-off date to study the effect of countermasures on coding tasks.
Nonetheless, owing to the distinctive features of code tasks, as we mentioned in Section II-A1, different conclusions from the NLP research concerning input/output-only contamination are obtained. Focusing on the domain of code intelligence, most studies only consider sample-level contamination [21, 23, 22], i.e., the fourth setting in this paper. As such, the effects of other fine-grained contamination issues remain unanswered questions, which is a striking gap this paper fills.
VIII Conclusion
@@
This paper presents a comprehensive empirical study to reveal the impact of different data contamination scenarios on code intelligence. The whole experiment covers diverse PLMs/LLMs, and code tasks, uncovering different conclusions that were released before in the NLP field, and challenging the conventional belief that contamination inevitably induces performance overestimation in the code intelligence field, manifesting the necessity of our research. Future work could further explore the effects of large-scale contamination and validate the generality of the findings across tasks and PLs.

### v3

@@ 6.1. Code Intelligence Tasks on PLMs and LLMs
Recent years have witnessed extensive research on code intelligence tasks based on PLMs/LLMs, such as code generation (Li et al., 2022; Le et al., 2022), code summarization (Ahmed and Devanbu, 2022; Geng et al., 2024), and code translation (Yang et al., 2024; Xue et al., 2025). Early efforts, such as CodeBERT (Feng et al., 2020) and GraphCodeBERT (Guo et al., 2020), jointly modeled NL and PL, enhancing semantic understanding of code through contrastive learning and structural awareness. Subsequently, PLBART (Ahmad et al., 2021), based on an encoder-decoder architecture, demonstrated strong transfer capabilities. CodeT5 (Wang et al., 2021) further pre-trained on T5 (Raffel et al., 2020), achieving state-of-the-art performance on a variety of code understanding and generation tasks. On the LLM front, Codex (Chen et al., 2021) and CodeGen (Nijkamp et al., 2022) have shown remarkable capabilities in multi-turn reasoning and complex programming scenarios. More recent works, such as StarCoder (Li et al., 2023) and WizardCoder (Luo et al., 2023) explore the upper performance bounds of open-source LLMs in code intelligence, emphasizing the importance of model scale, instruction tuning, and multilingual training for improving code comprehension and generation abilities.
6.2. Studies On Data Contamination
@@
Data contamination studies have been investigated for years, many are dedicated to contamination detection. For instance, Golchin et al. (Golchin and Surdeanu, 2024) proposed both instance-level and partition-level contamination detection, while Dong et al. (Dong et al., 2024) proposed CDD, using output distributions to detect contamination in LLMs on both code and NLP tasks. Deng et al. (Deng et al., 2024), instead, proposed TS-Guessing to detect contamination by examining whether LLMs can produce masked wrong answers. As for research exploring contamination effects, early research was conducted in the NLP field. For example, Magar et al. (Magar and Schwartz, 2022) explored to what extent BERT exploits or just memorises the contaminated data for downstream tasks. Jiang et al. (Jiang et al., 2024b) explored the input-only/output-only contamination effects on GPT-2. Subsequently, contamination has also been studied in the code intelligence field, Riddell et al. (Riddell et al., 2024) and Ramos et al. (Ramos et al., 2024) quantified the contamination effects on LLM-based code generation and bug fixing, respectively. Cao et al. (Cao et al., 2024) extracted data before and after LLMs’ cut-off date to study the effect of countermasures on coding tasks. Nonetheless, owing to the distinctive features of code tasks, as we mentioned in Section 2.1.1, different conclusions from the NLP research concerning input/output-only contamination are obtained. Focusing on the domain of code intelligence, most studies only consider sample-level contamination (Cao et al., 2024; Riddell et al., 2024; Ramos et al., 2024), i.e., the fourth setting in this paper. As such, the effects of other fine-grained contamination issues remain unanswered questions, which is a striking gap this paper fills.
7. Conclusion and Future Work
@@
This paper presents a fine-grained empirical study to reveal the impact of different data contamination scenarios on code intelligence. The whole experiment covers diverse PLMs/LLMs, PLs, and code tasks, uncovering different conclusions that were released before in the NLP field, and challenging the conventional belief that contamination inevitably induces performance overestimation in the code intelligence field, manifesting the necessity of our research. Future work could further explore the effects of larger-scale contamination and their influence on PLM/LLMs with post-training.